If the kinematic viscosity of water at S.TP is 1:0038 x 10 m²/a and the mass density is 998.23 kg/m³, calculate its absolute viscosity

1
Expert's answer
2022-02-16T14:54:02-0500

Absolute viscosity=(998.23"\frac{kg}{m^3}") /(1:0038"\times" 10m2/a)


Convert 998.23"\frac{kg}{m^3}" to g/mm3= "\frac{998.23}{1000}"= 0.99823"\frac{g}{mm^3}"



1:0038"\times" 10m2/a= "\frac{10m^2}{38a}" = "\frac{5m^2}{19a}"


Absolute viscosity= (0.99823"\frac{g}{mm^3}")/("\frac{5m^2}{19a}" )


Absolute viscosity= 0.26269Pa.S